# HabiCrowd: A High Performance Simulator for Crowd-Aware Visual Navigation

### v1

@@ Abstract
Visual navigation, a foundational aspect of Embodied AI (E-AI), has been significantly studied in the past few years. While many 3D simulators have been introduced to support visual navigation tasks, scarcely works have been directed towards combining human dynamics, creating the gap between simulation and real-world applications. Furthermore, current 3D simulators incorporating human dynamics have several limitations, particularly in terms of computational efficiency, which is a promise of E-AI simulators. To overcome these shortcomings, we introduce HabiCrowd, the first standard benchmark for crowd-aware visual navigation that integrates a crowd dynamics model with diverse human settings into photorealistic environments. Empirical evaluations demonstrate that our proposed human dynamics model achieves state-of-the-art performance in collision avoidance, while exhibiting superior computational efficiency compared to its counterparts. We leverage HabiCrowd to conduct several comprehensive studies on crowd-aware visual navigation tasks and human-robot interactions. The source code and data can be found at https://habicrowd.github.io/.
1 Introduction
Embodied AI (E-AI), the intersection between computer vision, machine learning, and robotics, has gained interest amongst scientists in the past few years ehsani2021manipulathor . Training E-AI agents that can perceive, reason, and interact with the environment offers significant potential for sim2real applications srivastava2022behavior . The progress in E-AI research has been rapidly accelerated yadav2023habitat thanks to the developments of fast and high-fidelity 3D simulators wijmans2019dd that enable large-scale computational parallelization savva2019habitat . Among all E-AI studies, one of the most fundamental problems is visual navigation gupta2017cognitive , which involves training an agent to navigate to a given goal using perception from the sensory inputs nguyen2019vision . While many benchmarks have been proposed to address the visual navigation tasks in computer vision wijmans2019dd ; nguyen2019help ; chaplot2020object ; bansal2020combining ; chen2020soundspaces , most of them assume that navigating environments are static and do not consider human dynamics. Conversely, agents typically interact within dynamic environments, particularly those featuring humans whose positions constantly change yi2023mime . If robots are to efficiently assist humans in routine duties, they must possess the capability to navigate effectively in crowded and dynamic environments monaci2022dipcan . Therefore, human dynamics is an essential factor that needs to be thoroughly examined when developing E-AI benchmarks bonin2008visual . In this paper, we explore the tasks of visual navigation for an agent in the presence of human dynamics, aiming to bridge the gap between simulation and real-life scenarios.
[…]
[width=]cvpr/figures/chapter1/sample/intro_figure.pdf
@@ 1 Introduction
The visual navigation task with the inclusion of human dynamics is widely referred to as the crowd-aware visual navigation task in the robotics community chen2019crowd . This task is challenging for several reasons. Firstly, to complete the navigation task, the agent needs to learn the patterns of human actions kretzschmar2016socially , which are complex and unstructured dugas2022navdreams . Furthermore, it is infeasible to predict the motion dynamics in crowded habitats as the motion of each person is typically affected by their neighbours alahi2017learning . Most prior works on crowd-aware navigation works in robotics usually consider training the agent in 2D simulators chen2019crowd ; borenstein1989real ; kretzschmar2016socially ; alahi2016social ; chen2017socially ; nguyen2022deep ; ferrer2013robot ; alahi2017learning ; trautman2020real ; nguyen2020autonomous ; liu2021decentralized ; bennewitz2005learning . The assumption of using 2D simulators is impractical dugas2022navdreams given that real-life robots primarily perceive human activities through visual sensors rather than employing mathematical social force models yadav2023habitat to address the task.
@@
Recently, a few attempts to add dynamic humans to 3D photorealistic simulators 2021igibsonchallenge ; makoviychuk2021isaac are shown to have limitations regarding the human dynamics model, as well as the available number of human models and scenes. In addition, one of the most critical drawbacks of recent 3D simulators integrating human dynamics lies in their computational efficiency. While E-AI simulators strive to facilitate rapid rendering and parallelization, prior simulators have proved to be inefficient in terms of computation yokoyama2022benchmarking ; makoviychuk2021isaac . In general, the lack of 3D simulators with well-designed human dynamics makes the developed methods less applicable and expresses the necessity of establishing a standardized 3D benchmark for simulators that integrate human behaviors.
@@
Motivated by these shortcomings, in this work, we introduce HabiCrowd, a new benchmark for the crowd-aware visual navigation task. HabiCrowd includes comprehensive social force-based human dynamics with diverse density settings operating in a large number of high-fidelity scenes from the Habitat-Matterport 3D dataset (HM3D) ramakrishnan2021hm3d . To assess the performance of HabiCrowd, several experiments are conducted, including benchmarking the human dynamics model, rendering speed, and memory utilization. Experimental results reveal that HabiCrowd achieves collision-free while running significantly faster than the state-of-the-art human dynamics model. Additionally, our HabiCrowd also exhibits a remarkable advantage in rendering speed and memory utilization compared to other related benchmarks. Upon HabiCrowd, we propose a new metric for crowd-aware navigation and benchmark two visual navigation tasks. The outcomes highlight the importance of human dynamics in navigating procedures. We also demonstrate that our new simulator enables various studies regarding human density and human-robot interactions. Our contributions are summarized as follows:
@@
  • We present HabiCrowd, a new dataset and benchmark for crowd-aware visual navigation. Experiments show that our simulator achieves a remarkable level of collision avoidance while surpassing other benchmarks in terms of human diversity and computational utilization.
@@
  • We benchmark navigation tasks on HabiCrowd and illustrate that our proposed simulator offers several in-depth analyses, such as studies of the impact of human density, human-robot avoidance, and human recognition on crowd-aware visual navigation tasks.
2 Related Works
@@
Visual Navigation. Traditionally, roboticists have addressed visual navigation tasks by utilizing control methods such as Model Predictive Control (MPC) rawlings2017model ; hirose2019deep . Several following works have been proposed to improve control-based methods sauvee2006image ; calli2017vision ; allibert2008real ; li2015vision ; zimmermann2023twostep ; whyte2006simultaneous . The idea of control methods has been conserved by utilizing the advantage of deep learning richter2018bayesian ; drews2017aggressive ; kaufmann2018deep ; bansal2020combining ; gupta2017cognitive ; finn2017deep ; hirose2019deep . Nevertheless, control-based methods demand a dynamics model and are often saturated when the number of ambiguities and the prediction horizon grows lucia2020stability . Recently, achievements in reinforcement learning (RL) mnih2015human have brought about a significant shift in visual navigation yang2018visual . Over the years, deep RL approaches have emerged as a widely adopted solution for addressing visual navigation tasks wijmans2019dd ; banino2018vector ; ma2019using ; ramakrishnan2022poni ; jianu2022cathsim ; yadav2023ovrl ; khandelwal2022simple ; ramrakhya2022habitat , and many RL-based methods have achieved state-of-the-art results in visual navigation tasks such as point-goal navigation wijmans2019dd , object-goal navigation ramrakhya2023pirlnav , image-goal navigation yadav2023ovrl .
@@
Crowd-aware Navigation. Navigation in crowded circumstances has significantly been examined by the robotics community chen2019crowd ; however, most works only employ 2D simulators or small-scale experiment setups kruse2013human . Earlier methods in crowd-aware navigation are primarily based on predicting the trajectories of human entities borenstein1989real ; bennewitz2005learning ; ferrer2013robot As the number of humans grows, the environment becomes exponentially denser; consequently, these trajectory-based approaches often suffer from the well-known freezing robot problem trautman2010unfreezing . Learning-based methods have been introduced to overcome the freezing robot issue by automatically determining optimal waypoints liu2021decentralized . Following this idea, RL frameworks are widely adopted to perceive human behaviors and interactions in a latent mechanism chen2019crowd ; kretzschmar2016socially ; alahi2016social ; chen2017socially ; alahi2017learning ; trautman2020real ; lo2019perception . However, RL methods based on 2D simulators often suppose that the dynamics of all human entities are explicit and well-determined while the real-life interactions and human dynamics are much more arbitrary and complex trautman2020real . To address this limitation, we study crowd-aware navigation in 3D settings, wherein the agent can only perceive the environment through egocentric inputs instead of 2D floor plans with known human dynamics.
@@
3D Simulators for Visual Navigation. There have been a large number of simulators for the tasks of visual navigation srivastava2022behavior ; moller2021survey . Gazebo, introduced in koenig2004design , is one of the first steps towards virtualizing the real world. Followed by Gazebo, several simulators have been established to study daily-conventional behaviors of robots, such as UnrealCV qiu2017unrealcv , AI2-THOR kolve2017ai2 , Gibson xia2018gibson , Habitat savva2019habitat ; szot2021habitat . Although the mentioned simulators provide opportunities to study interactions with humans, none of them have yet included human dynamics until the establishment of iGibson-Social Navigation (iGibson-SN) 2021igibsonchallenge ; 2022igibsonchallenge . More recently, Isaac Sim makoviychuk2021isaac is also established with the inclusion of virtual humans to the scenes. Nevertheless, the main drawbacks of both iGibson-SN and Isaac Sim are they have a limited number of scenes and human models. They also lack a measurement for crowdedness, which is an important factor in crowd-aware navigation problems  karamouzas2014universal . With that inspiration, we propose a human dynamics model that is carefully designed and included in Habitat 2.0 szot2021habitat to create a standard and diverse crowd-aware visual navigation benchmark that closely reflects real-world conditions. Table 1 indicates the main differences between our HabiCrowd, iGibson Social Navigation (iGibson-SN), and Isaac Sim simulator. The comparisons highlight that our simulator offers a greater range of navigation settings, scenes, and human models.
[…]
Table 1: Comparison with related simulators. We use the default settings nvidia_omniverse of Isaac Sim to benchmark scenes and virtual humans.
3 The HabiCrowd Simulator
@@
In this section, we present HabiCrowd, a crowd-aware visual navigation simulator utilizing Habitat 2.0 szot2021habitat as the foundational simulator. The motivation behind this choice comes from the diverse environment settings and photorealistic scenes offered by Habitat 2.0 yadav2023habitat . By integrating human dynamics into this simulator, HabiCrowd could further bridge the gap between simulation and real-life scenarios. Furthermore, Habitat 2.0 also enables parallel computing utilization deitke2022retrospectives , which could significantly facilitate the training procedure. We specify the configurations, trajectories, and controls of human dynamics of HabiCrowd.
3.1 Human Dynamics Model
@@
We design a continuous and force-based human dynamics model, which progresses concurrently with the agent’s navigation, based on the universal power law (UPL) dynamics model proposed by karamouzas2014universal . Previously, UPL lacks the inclusion of torque components to control the facing direction of humans. To address this limitation, we further incorporate the torque dynamics model into the simulator. Consequently, our improved dynamics model is denoted as UPL++.
@@
Assume the agent is navigating in a scene $\Lambda$ with the set of obstacles $W_{\Lambda}$ and $n$ humans. We represent each human by $\mathbf{P}_{i}=\left(\mathbf{x}_{i},\mathbf{v}_{i},\hat{\mathbf{e}}_{i},v_{i},\psi_{i},\psi_{i}^{0},\omega_{i},\omega_{i}^{0},r_{i},m_{i}\right)$ and consider their movements as particle kinematics on the 2D floor $\mathbf{F}$ of $\Lambda$. Table 3 explains the notations in this section. Each $\mathbf{P}_{i}$ will try to navigate to a destination $\mathbf{d}_{i}$. The model of $\mathbf{P}_{i}$ is depicted in Figure 2.
@@
Force Model. We utilize a force model for human entities to orient them to their destinations while minimizing collisions. The force model of $\mathbf{P}_{i}$ projecting on $\mathbf{F}$ can be written as:
@@
The fluctuation force $\zeta_{i}=\zeta_{i}^{\prime}\vec{n}\left(\varphi_{i}\right)$ is included to break the symmetry of all humans, where $\zeta_{i}^{\prime}\sim\mathcal{N}(0,\sigma_{\zeta}^{2})$, $\varphi_{i}\sim\mathcal{U}(0,2\pi)$, and $\vec{n}\left(\varphi_{i}\right)$ is the unit vector rotated by $\varphi_{i}$ about the origin of $\mathbf{F}$.
[…]
Not. Description
$r_{i}$ Particle’s radius
$\hat{\mathbf{e}}_{i}$ Desired direction
$\mathbf{x}_{i}$ Human coordinate
$\mathbf{v}_{i}$ Linear velocity
$v_{i}$ Desired velocity
$m_{i}$ Human mass
$\omega_{i}$ Current angular velocity
$\omega_{i}^{0}$ Desired angular velocity
@@
current orientation
$\psi_{i}^{0}$ Angular coordinate of the
@@
Figure 3: Notations of human dynamics.
The contact force $\mathbf{f}_{i,*}^{\text{con}}$, which is a consequence of a collision between other entities like objects/humans, indicates the corresponding physical reaction. In the former notation, *** is a placeholder indicating either an obstacle $w$ or another human index $j$. We rely on the core physic engine of Habitat 2.0, PyBullet coumans2019 to model the contact force.
@@
Finally, we establish the social force model between $\mathbf{P}_{i}$ and $\mathbf{P}_{j}$ to prolong their time-to-collision based on their linearly-estimated positions. Let $\tau$ be the time-to-collision between two entities and $\tilde{\mathbf{x}},\tilde{\mathbf{v}}$ be their relative position and velocity, respectively. Inspired by karamouzas2014universal , we define the interaction energy between $\mathbf{P}_{i}$ and $\mathbf{P}_{j}$, as follows:
@@
where $a=\|\tilde{\mathbf{v}}\|_{2}^{2};b=-\tilde{\mathbf{x}}^{\mathsf{T}}\tilde{\mathbf{v}};c=\|\tilde{\mathbf{x}}\|_{2}^{2}-(r_{i}+r_{j})^{2}$. Following karamouzas2014universal , we calculate the desired social force as $\mathbf{f}^{\text{soc}}_{i,j}(\tilde{\mathbf{x}})=-\nabla_{\tilde{\mathbf{x}}}E(\tau)$.
@@
Computational Complexity. We determine the computational complexity for the $i$-th virtual human. It can be implied from Eqs. (1) and (4) that the computation complexity for $\mathbf{f}_{i}$ and $\mathcal{M}_{i}$ is both $O(n)$. This is because, apart from $\mathbf{f}^{\text{soc}}_{i,j}$ and $\mathbf{f}^{\text{con}}_{i,j}$, which require $O(n)$ computations, the remaining components only need constant computations.
We note that our UPL++ dynamics model and ORCA van2011reciprocal that is used in iGibson-SN 2021igibsonchallenge share the same computational complexity of $O(n)$. However, the major difference between our UPL++ and ORCA is the approach used to determine the dynamic terms, $\mathbf{f}_{i}$ and $\mathcal{M}_{i}$. While UPL++ directly computes these terms, ORCA employs an indirect approach by solving an LP problem. Empirical results in Section 4.1 confirm that the constant factor associated with UPL++ is considerably smaller than ORCA’s, indicating a computational efficiency advantage in favor of UPL++.
Figure 4: Dataset construction pipeline.
[…]
Figure 4: Dataset construction pipeline.
Figure 5: Human statistics.
3.2 Dataset Construction
Figure 5 illustrates the pipeline to build HabiCrowd scenes. We begin with initializing parameters for each virtual human $\mathbf{P}_{i}$ and their destination $\mathbf{d}_{i}$. Next, we construct virtual humans using the MakeHuman briceno2018makehuman . Then the trajectories of humans are configured by using the dynamics model in Section 3.1 and are added to the scenes of HM3D dataset ramakrishnan2021hm3d to create the final scenes. Note that when a human entity reaches its destination, we navigate it back to its original position and repeat this loop to maintain human dynamics during the agent’s navigation.
Dataset Statistics. We utilize $56$ scenes from HM3D ramakrishnan2021hm3d and $40$ human entities. The train/val/test splits are $36/8/12$. HabiCrowd’s virtual humans have a gender ratio of $/1$ 1 and ages range from $15$ to $44$ (Figure 5). The human density and navigation area distributions are depicted in Figure 7. Our simulator provides a broad spectrum of navigation settings including different levels of human density, spanning from $0.1$ to $0.5$, as well as navigation areas ranging from $10\text{m}^{2}$ to $300\text{m}^{2}$. The average human density of HabiCrowd is $0.189$ humans per $\text{m}^{2}$, which is close to $0.143$, the real crowd density in the real world deloitte2010employment . Our simulator also has scenes with higher density (for example, 0.45) to make the problem more challenging.
Figure 6: Navigating statistics.
[…]

### v2

@@ Abstract
Visual navigation, a foundational aspect of Embodied AI (E-AI) and robotics has been extensively studied in the past few years. While many 3D simulators have been introduced for the visual navigation tasks, scarcely works have combined human dynamics, creating the gap between simulation and real-world applications. Furthermore, current 3D simulators incorporating human dynamics have several limitations, particularly in terms of computational efficiency, which is a promise of modern simulators. To overcome these issues, we introduce HabiCrowd, the new standard benchmark for crowd-aware visual navigation that includes a crowd dynamics model with diverse human settings into photorealistic environments. Empirical evaluations demonstrate that our proposed human dynamics model achieves state-of-the-art performance in collision avoidance while exhibiting superior computational efficiency compared to its counterparts. We leverage HabiCrowd to conduct several comprehensive studies on crowd-aware visual navigation tasks and human-robot interactions. The source code and data can be found at https://habicrowd.github.io/.
I Introduction
Embodied AI (E-AI), the intersection between computer vision, machine learning, and robotics, has gained interest amongst scientists in the past few years [1]. Training E-AI agents that can perceive, reason, and interact with the environment offers significant potential for sim2real applications [2]. The progress in E-AI research has been rapidly accelerated [3] thanks to the developments of fast and high-fidelity 3D simulators [4] that enable large-scale computational parallelization [5]. Among all E-AI studies, one of the most fundamental problems is visual navigation [6], which involves training an agent to navigate to a given goal using perception from the sensory inputs [7]. While many benchmarks have been proposed to address the visual navigation tasks [5, 8], most of them assume that navigating environments are static and do not consider human dynamics. Conversely, agents typically interact within dynamic environments, particularly those featuring humans whose positions constantly change [9]. If robots are to efficiently assist humans in routine duties, they must possess the capability to navigate effectively in crowded and dynamic environments [10]. Therefore, human dynamics is an essential factor that needs to be thoroughly examined when developing E-AI benchmarks [11]. This paper examines visual navigation in the presence of human dynamics to bridge the gap between simulation and real-life scenarios.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ I Introduction
The visual navigation task with the inclusion of human dynamics is widely referred to as the crowd-aware visual navigation task in the robotics community [12]. This task is challenging because the agent needs to learn the patterns of human actions [13], which are complex and unstructured [14]. Furthermore, it is infeasible to predict the motion dynamics in crowded habitats as the motion of each person is typically affected by their neighbours [15]. Most prior works train agents using 2D simulators [16, 9], which is impractical since robots perceive human activities via visual sensors rather than mathematical social force models [14].
@@
Recent attempts to add dynamic humans to 3D photorealistic simulators [17, 18] have shown limitations in human dynamics models and world scale. For instance, the limited number of humans employed in [17] can hinder generalization in human recognition [15]. Additionally, one of the most critical drawbacks of recent 3D simulators integrating human dynamics lies in their computational efficiency [19], even though E-AI simulators strive to facilitate rapid rendering and parallelization [4]. In general, the lack of 3D simulators with well-designed human dynamics reduces the applicability of developed methods, highlighting the need for a standardized 3D benchmark that integrates human behaviors.
@@
Motivated by these shortcomings, we introduce HabiCrowd, a new benchmark for the crowd-aware visual navigation task. HabiCrowd includes comprehensive social force-based human dynamics with diverse density settings operating across numerous high-fidelity scenes from the Habitat-Matterport 3D dataset (HM3D) [7]. To assess the performance of HabiCrowd, several experiments are conducted, including benchmarking the human dynamics model, rendering speed, and memory utilization. Experimental results reveal that HabiCrowd achieves collision-free while running significantly faster than the state-of-the-art human dynamics model. Additionally, our HabiCrowd also exhibits a remarkable advantage in rendering speed and memory utilization compared to other related benchmarks. Upon HabiCrowd, we propose a new metric for crowd-aware navigation and benchmark two visual navigation tasks. The outcomes highlight the importance of human dynamics in navigating procedures. We also demonstrate that our new simulator enables various studies regarding human density and human-robot interactions. Our contributions are summarized as follows:
@@
  • We present HabiCrowd, a new dataset and benchmark for crowd-aware visual navigation. Experiments show that our simulator excels in terms of performance, human diversity and computational efficiency.
@@
  • We comprehensively benchmark navigation tasks using HabiCrowd to showcase that our simulator offers detailed analyses, such as studies on human density and human-robot avoidance in crowd-aware environments.
II Related Works
@@
Visual Navigation. Traditionally, roboticists have addressed visual navigation tasks by utilizing control methods such as Model Predictive Control (MPC) [20]. Several following works have been proposed to improve control-based methods [21]. The idea of control methods has been conserved by utilizing the advantage of deep learning [20]. Nevertheless, control-based methods demand a dynamics model and are often saturated when the number of ambiguities and the prediction horizon grows [22]. Recently, achievements in reinforcement learning (RL) [23] have brought about a significant shift in visual navigation [24]. Over the years, deep RL approaches have emerged as a widely adopted solution for addressing visual navigation tasks [1, 25], and many RL-based methods have achieved state-of-the-art results in visual navigation tasks such as point-goal navigation [4], object-goal navigation [26], image-goal navigation [1].
@@
Crowd-aware Navigation. Navigation in crowded circumstances has significantly been examined by the robotics community [12]. Earlier methods in crowd-aware navigation are primarily based on predicting the trajectories of human entities [27]. As the number of humans grows, the environment becomes exponentially denser; consequently, these trajectory-based approaches often suffer from the well-known freezing robot problem [28]. Learning-based methods have been introduced to overcome the freezing robot issue by automatically determining optimal waypoints [9]. Following this idea, RL frameworks are widely adopted to perceive human behaviors and interactions in a latent mechanism [12, 29]. However, RL methods based on 2D simulators often suppose that the dynamics of all human entities are explicit and well-determined while the real-life interactions and human dynamics are much more arbitrary and complex [29]. To address this limitation, we study crowd-aware navigation in 3D settings, wherein the agent can only perceive the environment through egocentric inputs instead of 2D floor plans with known human dynamics.
@@
3D Simulators for Visual Navigation. There have been a large number of simulators for the tasks of visual navigation [30]. Gazebo, introduced in [31], is one of the first steps towards virtualizing the real world. Followed by Gazebo, several simulators have been established to study daily-conventional behaviors of robots, such as UnrealCV [32], AI2-THOR [33], Gibson [34], Habitat [35]. Although the mentioned simulators provide opportunities to study interactions with humans, none of them have yet included human dynamics until the establishment of iGibson-Social Navigation (iGibson-SN) [17]. More recently, Isaac Sim [18] is also established with the inclusion of virtual humans to the scenes. Nevertheless, the main drawbacks of both iGibson-SN and Isaac Sim are they have a limited number of scenes and human models. They also lack a measurement for crowdedness, which is an important factor in crowd-aware navigation problems  [36]. With that inspiration, we propose a human dynamics model that is carefully designed and included in Habitat 2.0 [35] to create a standard and diverse crowd-aware visual navigation benchmark that closely reflects real-world conditions. Table I indicates the main differences between our HabiCrowd, iGibson Social Navigation (iGibson-SN), and Isaac Sim simulator. The comparisons underscore the substantial advantages of our simulator, which offers a significantly greater number of scenes and a wider range of navigation settings, environments, and human models.
TABLE I: Simulator Comparison.
Dynamics model Navigation settings Num. scenes Environment type Num. humans
iGibson-SN [17] ORCA [37] Point-goal 15 residence 3
Isaac Sim [18] Point-goal 7 office, depot traffic, hospital 7
HabiCrowd (ours) UPL++ Point-goal Object-goal Image-goal 480 residence, gym, office, studio, club, restaurant 40
III The HabiCrowd Simulator
@@
In this section, we present HabiCrowd, a crowd-aware visual navigation simulator utilizing Habitat 2.0 [35] as the foundational simulator. The motivation behind this choice comes from the diverse environment settings and photorealistic scenes offered by Habitat 2.0 [3]. By integrating human dynamics into this simulator, HabiCrowd could further bridge the gap between simulation and real-life scenarios. Furthermore, Habitat 2.0 also enables parallel computing utilization [38], which could significantly facilitate the training procedure. We specify the configurations, trajectories, and controls of human dynamics of HabiCrowd.
III-A Human Dynamics Model
@@
We design a continuous and force-based human dynamics model, which progresses concurrently with the agent’s navigation, based on the universal power law (UPL) dynamics model proposed by [36]. Previously, UPL lacks the inclusion of torque components to control the facing direction of humans. To address this limitation, we further incorporate the torque dynamics model into the simulator. Consequently, our improved dynamics model is denoted as UPL++.
@@
Assume the agent is navigating in a scene $\Lambda$ with the set of obstacles $W_{\Lambda}$ and $n$ humans. Table II explains the notations in this section. We represent each human by $\mathbf{P}_{i}=\left(\mathbf{x}_{i},\mathbf{v}_{i},\hat{\mathbf{e}}_{i},v_{i},\psi_{i},\psi_{i}^{0},\omega_{i},\omega_{i}^{0},r_{i},m_{i}\right)$ and consider their movements as particle kinematics on the 2D floor $\mathbf{F}$ of $\Lambda$. Each $\mathbf{P}_{i}$ will try to navigate to a destination $\mathbf{d}_{i}$. The model of $\mathbf{P}_{i}$ is depicted in Figure 2. We denote $\mathbf{f}_{i}$ and $\mathcal{M}_{i}$ as the applied force and torque to each $\mathbf{P}_{i}$, with the subscripts ”adj,” ”soc,” and ”con” corresponding ”adjective,” ”social,” and ”contact”, respectively.
@@
Force Model. We use a force model for human entities to guide them to their destinations while minimizing collisions. The force model of $\mathbf{P}_{i}$ projecting on $\mathbf{F}$ can be written as:
@@
The fluctuation force $\zeta_{i}=\zeta_{i}^{\prime}\mathbf{n}\left(\varphi_{i}\right)$ is included to break the symmetry of all humans, where $\zeta_{i}^{\prime}\sim\mathcal{N}(0,\sigma_{\zeta}^{2})$, $\varphi_{i}\sim\mathcal{U}(0,2\pi)$, and $\mathbf{n}\left(\varphi_{i}\right)$ is the unit vector rotated by $\varphi_{i}$ about the origin of $\mathbf{F}$.
Figure 2: Human dynamics model. We design a model to shape $\mathbf{P}_{i}$ towards $\mathbf{d}_{i}$ while avoiding collisions with $\mathbf{P}_{j}$. The human dynamics includes a force model and a torque model.
TABLE II: Human Dynamics Notations.
Not. Description Not. Description
$r_{i}$ Particle’s radius $\hat{\mathbf{e}}_{i}$ Desired direction
$\mathbf{x}_{i}$ Human coordinate $\mathbf{v}_{i}$ Linear velocity
$v_{i}$ Desired velocity $m_{i}$ Human mass
$\omega_{i}$ Current angular velocity $\omega_{i}^{0}$ Desired angular velocity
@@
current orientation $\psi_{i}^{0}$ Angular coordinate of the
@@
Figure 3: Dataset construction pipeline.
Figure 4: Human statistics.
The contact force $\mathbf{f}_{i,*}^{\text{con}}$, which is a consequence of a collision between other entities like objects/humans, indicates the corresponding physical reaction. In the former notation, * is a placeholder indicating either an obstacle $w$ or another human index $j$. We rely on the core physic engine of Habitat 2.0, PyBullet [39] to model the contact force.
@@
Finally, we establish the social force model between $\mathbf{P}_{i}$ and $\mathbf{P}_{j}$ to prolong their time-to-collision based on their linearly-estimated positions. Let $\tau$ be the time-to-collision between two entities and $\tilde{\mathbf{x}},\tilde{\mathbf{v}}$ be their relative position and velocity, respectively. Inspired by [36], we define the interaction energy between $\mathbf{P}_{i}$ and $\mathbf{P}_{j}$, as follows:
@@
where $a=\|\tilde{\mathbf{v}}\|_{2}^{2};b=-\tilde{\mathbf{x}}^{\mathsf{T}}\tilde{\mathbf{v}};c=\|\tilde{\mathbf{x}}\|_{2}^{2}-(r_{i}+r_{j})^{2}$. Following [36], we calculate the desired social force as $\mathbf{f}^{\text{soc}}_{i,j}(\tilde{\mathbf{x}})=-\nabla_{\tilde{\mathbf{x}}}E(\tau)$.
@@
Computational Complexity. It can be implied from Eqs. (1) and (4) that the computation complexity for $\mathbf{f}_{i}$ and $\mathcal{M}_{i}$ is both $O(n)$. This is because, apart from $\mathbf{f}^{\text{soc}}_{i,j}$ and $\mathbf{f}^{\text{con}}_{i,j}$, which require $O(n)$ computations, the remaining components only need constant computations.
We note that our UPL++ dynamics model and ORCA [37] that is used in iGibson-SN [17] share the same computational complexity of $O(n)$. However, the major difference between our UPL++ and ORCA is the approach used to determine the dynamic terms, $\mathbf{f}_{i}$ and $\mathcal{M}_{i}$. While UPL++ directly computes these terms, ORCA employs an indirect approach by solving an LP problem. Empirical results in Section IV-A confirm that the constant factor associated with UPL++ is considerably smaller than ORCA’s, indicating a computational efficiency advantage in favor of UPL++.
III-B Dataset Construction
Figure 4 illustrates the pipeline to build HabiCrowd scenes. We begin with initializing parameters for each virtual human $\mathbf{P}_{i}$ and their destination $\mathbf{d}_{i}$. Next, we construct virtual humans using the MakeHuman [40]. Then the trajectories of humans are configured by using the dynamics model in Section III-A and are added to the scenes of HM3D dataset [7] to create the final scenes. Note that when a human entity reaches its destination, we navigate it back to its original position and repeat this loop to maintain human dynamics during the agent’s navigation.